OpenStudy (mendicant_bias):

(ODE) Alright, working through a BVP, legit not sure how to deal with these, lol...

OpenStudy (mendicant_bias):

http://i.imgur.com/cIikLdS.png

OpenStudy (mendicant_bias):

I'm not quite sure how you deal with BVP's differently or what makes them fundamentally different from an IVP.

OpenStudy (mendicant_bias):

"There is an important mathematical difference between these kinds of problems: if we have a second order differential equation with given initial values whether or not a unique solution exists depends only on the equation, not the initial conditions, while whether or not there exists a solution to the same equation, with given boundary conditions, depends upon the boundary conditions as well." http://mathhelpforum.com/differential-equations/180630-difference-between-ivp-bvp.html

OpenStudy (mendicant_bias):

Alright, so that doesn't really mean anything to me practically, but hey! lol. Food for thought.

OpenStudy (mendicant_bias):

"Difference between an Initial Value Problem(IVP) and a Boundary Value Problem(BVP): - IVP: all the values needed to solve the particular problem are specified at a single point. - BVP: all the values needed to solve the particular problem are specified at different points." Just putting this here to think about. Still not sure how to approach this problem, gonna go read about it for a sec.

OpenStudy (mendicant_bias):

Okay, it looks like even if the conditions of that theorem 4.1.1 ( http://i.imgur.com/sLoesmC.png) are satisfied, there may be one solution, many (infinitely many, possibly) solutions, or none at all.

ganeshie8 (ganeshie8):

IVP : y(0) = .., y(0) = .. BVP : y(0) = .., y(1) = ..

ganeshie8 (ganeshie8):

notice that in BVP the points given are at different times : x = 0 and x = 1

ganeshie8 (ganeshie8):

but the points given in IVP are at same times : x = 0 and x = 0

OpenStudy (mendicant_bias):

Yeah, I see that, I'm just wondering mechanically how to approach this.

ganeshie8 (ganeshie8):

you approach it exactly same as the IVP that you have solved earlier

ganeshie8 (ganeshie8):

get two equations and solve the constants

OpenStudy (mendicant_bias):

This has to be in some way fundamentally different from the first way we attempted to solve VP's, because it seems like you'll have to be simultaneously plugging in values for the same variable, but that makes no sense to me. Alright, I'll take a shot or something like that, lol.

OpenStudy (mendicant_bias):

Wait, I think I got it

OpenStudy (mendicant_bias):

\[c_{1}+c_{2}=y=0.\]\[c_{1}e+\frac{c_{2}}{e}=1\]

OpenStudy (mendicant_bias):

Alright, now just algebra.

ganeshie8 (ganeshie8):

yeah give it to wolfram :O

OpenStudy (mendicant_bias):

I can see how to solve this by substitution and stuff, I think I'm-lol

OpenStudy (mendicant_bias):

Okay, cool, makes sense. On to another problem type. Thank you for helping me through this stuff, it means a lot.

ganeshie8 (ganeshie8):

lol that kid is a psycho

OpenStudy (mendicant_bias):

kid? Wait, what kid, wot

OpenStudy (mendicant_bias):

(Lol what)

ganeshie8 (ganeshie8):

oh you're done with b and c ?

OpenStudy (mendicant_bias):

...............

OpenStudy (mendicant_bias):

See step 15. http://www.smbc-comics.com/comics/20130616.png

OpenStudy (mendicant_bias):

But yeah, taking a look at B, thank you for reminding me, in all seriousness.

OpenStudy (mendicant_bias):

(Alright, be right back, like, literally 3 minutes DON'T GO ANYWHERE :E)

ganeshie8 (ganeshie8):

sorry to remind you about more work lol but part b looks similar to part a and you can finish off part c somehow..

OpenStudy (mendicant_bias):

Alright, I'm back. Yup, gonna take a shot.

OpenStudy (mendicant_bias):

I'm not sure, but I think I might end up having to use the exponential definitions of cosh and sinh. Either way, setting up the general form now.

OpenStudy (mendicant_bias):

\[c_{3}\cosh(0)+c_{4}\sinh(0)=0;\]\[c_{3}\cosh(1)+c_{4}\sinh(1)=1.\]

OpenStudy (mendicant_bias):

Alright, erm.......lol, better go look for hyperbolic identities or properties, or otherwise bust out the exponential definitions of both.

OpenStudy (mendicant_bias):

Well, I'm not really familiar with the functions, but apparently

OpenStudy (mendicant_bias):

\[\cosh(0)=1, \ \sinh(0)=1.\]

OpenStudy (mendicant_bias):

\[c_{3}+c_{4}=0; c_{3}=-c_{4}.\]

OpenStudy (mendicant_bias):

\[-c_{4}\cosh(1)+c_{4}\sinh(1)=1; \ \ \ c_{4}(\sinh(1)-\cosh(1))=1.\]

OpenStudy (mendicant_bias):

Time to look for identities.

OpenStudy (mendicant_bias):

Aaaaaaand nothing of value was gained.

ganeshie8 (ganeshie8):

your first equation directly gives you \(c_3 = 0\)

OpenStudy (mendicant_bias):

(?)

ganeshie8 (ganeshie8):

\[c_{3}\cosh(0)+c_{4}\sinh(0)=0 \] \[c_{3}(1)+c_{4}(0)=0 \] \[c_3 = 0\] right ?

OpenStudy (mendicant_bias):

I plugged into google both cosh(0) and sinh(0) and got them to be identically one, not zero, lemme double check on that.

OpenStudy (mendicant_bias):

Nah, you're right, understood.

OpenStudy (mendicant_bias):

If \[c_{3}=0, \]then \[c_{4}\sinh(1)=1, c_{4}=\frac{1}{\sinh(1)}\]

OpenStudy (mendicant_bias):

Which is a really nasty value numerically.

ganeshie8 (ganeshie8):

not really, use the definition of hyperbolic functions

OpenStudy (mendicant_bias):

Alright, lemme look 'em up, one sec

ganeshie8 (ganeshie8):

\[\sinh x = \dfrac{e^x-e^{-x}}{2}\]

OpenStudy (mendicant_bias):

\[\cosh(x)=\frac{e^{x}+e^{-x}}{2}, \ \ \ \sinh(x)=\frac{e^{x}-e^{-x}}{2}\]

OpenStudy (mendicant_bias):

Oh, and since it's reciprocal

OpenStudy (mendicant_bias):

We have like, \[\rm csch(1)=c_{4}\]

OpenStudy (mendicant_bias):

\[\sinh(1)=\frac{e^{1}-e^{-1}}{2}=\frac{e}{2}-\frac{1}{2e}\]

OpenStudy (mendicant_bias):

So the reciprocal of the entire right side is our answer expressed by the definition.

OpenStudy (mendicant_bias):

Looks like we made mistakes with both part A and part B? http://i.imgur.com/OFQN6L9.png

ganeshie8 (ganeshie8):

\[y = 0\cosh x +\frac{1}{\sinh(1)} \sinh (x)\]

ganeshie8 (ganeshie8):

you need to plug the constants back to the given solution ^

OpenStudy (mendicant_bias):

Yeah, nevermind, p1 was good, and you're right, that's always the thing I'm comically worst at remembering to do, the easiest and last part.

OpenStudy (mendicant_bias):

Alright, well, cool, problem solved. Going to move on to some more stuff.

OpenStudy (mendicant_bias):

Exam in 13 hours.

ganeshie8 (ganeshie8):

Ohk dont waste time on part c, its just algebra..